# MLAAD: The Multi-Language Audio Anti-Spoofing Dataset

### v6

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems. To address this limitation, this paper presents the Multi-Language Audio Anti-Spoof Dataset (MLAAD), created using 91 TTS models, comprising 42 different architectures, to generate 420.7 hours of synthetic voice in 38 different languages.
@@
By publishing [^1] https://deepfake-total.com/mlaad MLAAD and making a trained model accessible via an interactive webserver [^2] https://deepfake-total.com/ , we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, and contributing to global efforts against audio spoofing and deepfakes.
@@ I Introduction
To address these issues, we introduce our MLAAD dataset, a large-scale corpus of state-of-the-art audio fakes encompassing 38 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 420.7 hours of synthesized speech, created by 91 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ II Related Work
MLAAD (ours) 38 91 175,000
TABLE I: Overview of Related Work on Datasets Used in Audio Anti-spoofing Research.
@@ II-C Datasets for Deepfake Detection
Several voice anti-spoofing datasets have been released for training these models, c.f. Table I. These datasets focus on enhancing weak model generalization capability by including a wide array of generation methods, codecs, noises, and data quality [12, 17, 18, 19, 22], as well as dealing with partial spoofs [16, 21, 18], real-world deepfake instances [6], and multimodal deepfakes combining fake audio and video [13, 58]. However, a gap remains in addressing the diversity of languages in spoofed utterances, with most datasets focusing predominantly on English [12, 16, 6, 17, 13] and Chinese [21, 22, 20, 18, 19]. Although WaveFake [17] incorporates both English and Japanese, datasets predominantly featuring English continue to dominate the field of anti-spoofing. To the best of our knowledge, there is no speech anti-spoofing dataset that comprehensively addresses multiple languages.
@@ III-A Synthesis Procedure
We introduce the MLAAD dataset, an extension of the M-AILABS Speech Dataset, which features audio clips in eight original languages: English, French, German, Italian, Polish, Russian, Spanish, and Ukrainian. All of these audio clips represent authentic human speech, taken from audiobooks or speeches and interviews of public figures, such as former German Chancellor Angela Merkel. We enrich this dataset with computer-generated audio clips in 38 languages.
@@
Third, we aggregate the results: information on the synthesized audio file, its language, and the transcript are stored in a meta.csv file within the corresponding folder. These files constitute the “fake” audio files when used for supervised learning, while the original baseline audio files are used as “bona-fide” training samples. We use 22.050 Hz, with a 16-bit depth, and store the audio in WAV format. The TTS models are derived from Coqui.ai and Hugging Face and encompass nearly all current state-of-the-art TTS models, c.f. Fig. 2. Additionally, for each of the original languages present in M-AILABS, we use Griffin Lim [60] as a model drop-in to re-synthesize 1000 randomly chosen audio files.
@@ III-B Formatting
TABLE II: Descriptors in the MLAAD dataset.
The outputs are organized into directories describing language and model type. Each directory contains the synthesized audio files and an accompanying “meta.csv” file, which encapsulates various data fields as indicated in  Table II. The “path” field specifies the location of the synthesized audio file, relative to the base path of MLAAD. The “original_file” field denotes the filepath of the baseline audio, relative to the base path of the M-AILABS dataset. The “language” field identifies the language of the audio file, while “is_original_language” indicates if the language was originally included in M-AILABS, or if the transcript has been translated. The “duration” field shows the length of the audio file in seconds, and “training_data” reveals the dataset used to train the TTS model. The “model_name” field provides a unique identifier for the model, such as “facebook/mms-tts-eng” for the Facebook Massively Multilingual Speech (MMS) TTS model [61], while “architecture” describes the TTS model architecture (such as VITS). Lastly, “transcript” represents the textual transcript of the audio file.
@@ III-C Statistics
The dataset encompasses 420.7 hours of synthesized speech across 38 languages, generated using 91 state-of-the-art TTS models, comprising 42 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2, along with Table V and Table VI, presents the architecture, dataset, and language distribution within MLAAD, offering an overview of the TTS architectures, training datasets, and languages that constitute the dataset. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single-speaker datasets [39, 43, 44, 62, 40], as well as larger, multi-language and multi-speaker databases [41, 42, 10]. For some of the models, the training database is either unknown, non-existent (e.g., Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [39] and audio recorded through community-driven initiatives [42].
@@ IV Evaluation
Model Train Data $\downarrow$ ASVspoof19 ASVspoof21-DF ASVspoof21-LA FakeOrReal InTheWild MLAAD Voc.v WaveFake
@@
MLAAD 70.8±0.9 52.7±6.1 52.6±6.2 50.5±3.3 54.3±4.9 - 86.7±29.4 97.2±3.3
TABLE III: Summary of training results across different datasets, measured by test accuracy. Each row represents a model trained on one dataset and evaluated on the remaining seven datasets, as described in Section IV-B. Highlights indicate the highest value per test dataset (i.e., per column). Averages are calculated from five separate trials, presented as mean ± standard deviation.
@@ IV-B Cross-dataset evaluation
These learning models are trained separately on four different datasets: MLAAD, ASVspoof19 [4], FakeOrReal (FoR) [14], and In-The-Wild (ITW) [6]. Evaluation is conducted on eight datasets, including the four used for training, as well as ASVspoof21-DF and ASVspoof21-LA [12], Voc.v [15], and WaveFake [17]. We do not use the latter four datasets for training because ASVspoof21-LA and DF are intended solely as test-sets [12]. Voc.v and WaveFake consist only of fake samples, precluding supervised training. Finally, we refrain from using ASVspoof21-LA to assess the cross-dataset performance of models trained on ASVspoof19, given the substantial similarity between the two datasets.
For every configuration, we conduct the experiment five times, reporting the mean accuracy along with the standard deviation in Table III. The results quantify a dataset’s ability to enable state-of-the-art models to generalize voice anti-spoofing to cross-dataset data, a key requirement for real-world applicability.
@@ IV-C Assessment of Synthesis Quality
The outcomes are detailed in Table IV. In many cases, the error rates observed between the synthesized audio and the original audio are comparable. This is true especially for languages present in the original M-AILABS dataset, as well as other central European languages such as Dutch, Czech or Finish. However, in cases where the alphabet significantly diverges from the ASCII standard, such as Hindi, Maltese, Swahili or Ukrainian, we observe a noticeable increase in error rates. Despite the TTS models (and possibly the ASR systems) supporting numerous languages, a substantial disparity in quality remains evident. Prominent languages like English, Spanish or German exhibit lower error rates, whereas others encounter more challenges in accurate processing. This variation underscores the importance of enhancing support for underrepresented languages, as the current performance disparity disproportionately affects individuals from these communities.
@@
TABLE IV: Quality of the MLAAD dataset, as measured in mean edit distance $\text{ED}_{s}$ (Fake) and $\text{ED}_{o}$ (Original), with standard deviation shown. Edit distance is computed between the original text (or translated version thereof) and the transcript of the corresponding audio file.
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 38 languages by 91 TTS models, composed of 42 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.

### v7

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems. To address this limitation, this paper presents the Multi-Language Audio Anti-Spoofing Dataset (MLAAD), version 7, created using 101 TTS models, comprising 52 different architectures, to generate 485.3 hours of synthetic voice in 40 different languages.
@@
By publishing [^1] https://deepfake-total.com/mlaad and making a trained model accessible via an interactive webserver [^2] https://deepfake-total.com/ , we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, and contributing to global efforts against audio spoofing and deepfakes.
@@ I Introduction
Release Date Hours Models Languages Distributed via
v1 17.01.2024
& 21 22 zip v2 28.02.2024 63.0 21 22 zip v3 16.04.2024 163.9 54 23 zip v4 24.09.2024 175.0 59 23 zip v5 02.11.2024 398.3 87 38 zip + Huggingface v6 26.04.2025 420.7 91 38 Huggingface v7 11.07.2025 485.3 101 40 Huggingface
TABLE I: Overview of MLAAD dataset versions, including release dates, total duration of speech data, number of TTS models, supported languages, and distribution method.
@@
To address these issues, we introduce and continuosly expand our MLAAD dataset (c.f. Section I), a large-scale corpus of state-of-the-art audio fakes encompassing 40 languages. This dataset, based on the M-AILABS Speech Dataset [10], includes 485.3 hours of synthesized speech, created by 101 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ II Related Work
MLAAD (ours) 40 101 201,000
TABLE II: Overview of Related Work on Datasets Used in Audio Anti-spoofing Research.
@@ II-C Datasets for Deepfake Detection
Several voice anti-spoofing datasets have been released for training these models, c.f. Table II. These datasets focus on enhancing weak model generalization capability by including a wide array of generation methods, codecs, noises, and data quality [12, 17, 18, 19, 22], as well as dealing with partial spoofs [16, 21, 18], real-world deepfake instances [6], and multimodal deepfakes combining fake audio and video [13, 58]. However, a gap remains in addressing the diversity of languages in spoofed utterances, with most datasets focusing predominantly on English [12, 16, 6, 17, 13] and Chinese [21, 22, 20, 18, 19]. Although WaveFake [17] incorporates both English and Japanese, datasets predominantly featuring English continue to dominate the field of anti-spoofing. To the best of our knowledge, there is no speech anti-spoofing dataset that comprehensively addresses multiple languages.
@@ III-A Synthesis Procedure
We introduce the MLAAD dataset, an extension of the M-AILABS Speech Dataset, which features audio clips in eight original languages: English, French, German, Italian, Polish, Russian, Spanish, and Ukrainian. All of these audio clips represent authentic human speech, taken from audiobooks or speeches and interviews of public figures, such as former German Chancellor Angela Merkel. We enrich this dataset with computer-generated audio clips in 40 languages.
@@
Third, we aggregate the results: information on the synthesized audio file, its language, and the transcript are stored in a meta.csv file within the corresponding folder. These files constitute the “fake” audio files when used for supervised learning, while the original baseline audio files are used as “bona-fide” training samples. We use 22,050 Hz, with a 16-bit depth, and store the audio in WAV format. The TTS models are derived from Coqui.ai and Hugging Face and encompass nearly all current state-of-the-art TTS models, c.f. Fig. 2. Additionally, for each of the original languages present in M-AILABS, we use Griffin Lim [60] as a model drop-in to re-synthesize 1000 randomly chosen audio files.
@@ III-B Formatting
TABLE III: Descriptors in the MLAAD dataset.
The outputs are organized into directories describing language and model type. Each directory contains the synthesized audio files and an accompanying “meta.csv” file, which encapsulates various data fields as indicated in  Table III. The “path” field specifies the location of the synthesized audio file, relative to the base path of MLAAD. The “original_file” field denotes the filepath of the baseline audio, relative to the base path of the M-AILABS dataset. The “language” field identifies the language of the audio file, while “is_original_language” indicates if the language was originally included in M-AILABS, or if the transcript has been translated. The “duration” field shows the length of the audio file in seconds, and “training_data” reveals the dataset used to train the TTS model. The “model_name” field provides a unique identifier for the model, such as “facebook/mms-tts-eng” for the Facebook Massively Multilingual Speech (MMS) TTS model [61], while “architecture” describes the TTS model architecture (such as VITS). Lastly, “transcript” represents the textual transcript of the audio file.
@@ III-C Statistics
The dataset encompasses 485.3 hours of synthesized speech across 40 languages, generated using 101 state-of-the-art TTS models, comprising 52 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2, along with Table VI and Table VII, presents the architecture, dataset, and language distribution within MLAAD, offering an overview of the TTS architectures, training datasets, and languages that constitute the dataset. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single-speaker datasets [39, 43, 44, 62, 40], as well as larger, multi-language and multi-speaker databases [41, 42, 10]. For some of the models, the training database is either unknown, non-existent (e.g., Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [39] and audio recorded through community-driven initiatives [42].
@@ IV Evaluation
Model Train Data $\downarrow$ ASVspoof19 ASVspoof21-DF ASVspoof21-LA FakeOrReal InTheWild MLAAD v1 Voc.v WaveFake
@@
MLAAD v1 70.8±0.9 52.7±6.1 52.6±6.2 50.5±3.3 54.3±4.9 - 86.7±29.4 97.2±3.3
TABLE IV: Summary of training results across different datasets, measured by test accuracy. Each row represents a model trained on one dataset and evaluated on the remaining seven datasets, as described in Section IV-B. Highlights indicate the highest value per test dataset (i.e., per column). Averages are calculated from five separate trials, presented as mean ± standard deviation.
@@ IV-B Cross-dataset evaluation
These learning models are trained separately on four different datasets: MLAAD v1, ASVspoof19 [4], FakeOrReal (FoR) [14], and In-The-Wild (ITW) [6]. Evaluation is conducted on eight datasets, including the four used for training, as well as ASVspoof21-DF and ASVspoof21-LA [12], Voc.v [15], and WaveFake [17]. We do not use the latter four datasets for training because ASVspoof21-LA and DF are intended solely as test-sets [12]. Voc.v and WaveFake consist only of fake samples, precluding supervised training. Finally, we refrain from using ASVspoof21-LA to assess the cross-dataset performance of models trained on ASVspoof19, given the substantial similarity between the two datasets.
For every configuration, we conduct the experiment five times, reporting the mean accuracy along with the standard deviation in Table IV. The results quantify a dataset’s ability to enable state-of-the-art models to generalize voice anti-spoofing to cross-dataset data, a key requirement for real-world applicability.
@@ IV-C Assessment of Synthesis Quality
The outcomes are detailed in Table V. In many cases, the error rates observed between the synthesized audio and the original audio are comparable. This is true especially for languages present in the original M-AILABS dataset, as well as other central European languages such as Dutch, Czech or Finish. However, in cases where the alphabet significantly diverges from the ASCII standard, such as Hindi, Maltese, Swahili or Ukrainian, we observe a noticeable increase in error rates. Despite the TTS models (and possibly the ASR systems) supporting numerous languages, a substantial disparity in quality remains evident. Prominent languages like English, Spanish or German exhibit lower error rates, whereas others encounter more challenges in accurate processing. This variation underscores the importance of enhancing support for underrepresented languages, as the current performance disparity disproportionately affects individuals from these communities.
@@
TABLE V: Quality of the MLAAD v1 dataset, as measured in mean edit distance $\text{ED}_{s}$ (Fake) and $\text{ED}_{o}$ (Original), with standard deviation shown. Edit distance is computed between the original text (or translated version thereof) and the transcript of the corresponding audio file.
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 40 languages by 101 TTS models, composed of 52 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.